OpenStudy (superkay):

5+4×2+6-2×2-1 Insert parenthesis so that it has a value of 19.

Awolflover1 (awolflover1):

\[\left(5+4\right)\times 2+6-2\times 2-1\]

OpenStudy (superkay):

Thanks

Awolflover1 (awolflover1):

you're welcome